OpenStudy (anonymous):

te^-t^2 chain rule Im not sure of the steps. thanks

OpenStudy (anonymous):

What do you want to do? Integrate that function?

OpenStudy (anonymous):

differentiate

OpenStudy (anonymous):

so, it's \[\frac{d(te^{-t^2})}{dt}\]

OpenStudy (anonymous):

well yes using the chain rule

OpenStudy (anonymous):

use product rule first. Can you?

OpenStudy (anonymous):

well I was trying to do that first but it didnt work

OpenStudy (anonymous):

\[f'(t)=e^{-t^2}+t\frac{de^{-t^2}}{dt}\] now we use chain rule to find \[\frac{de^{-t^2}}{dt}\]

OpenStudy (anonymous):

I was kidding, we can't use chain rule only here.